OpenStudy (anonymous):

f(x) = e^3 * ln(x) f'(x)=

OpenStudy (mimi_x3):

\[\large f(x) = e^3 * lnx \] \[\large f”(x) = e^3*\frac{1}{x} =>\frac{e^{3}}{x} \]

OpenStudy (anonymous):

e^3 is just a constant so....or just look at the answer.

OpenStudy (accessdenied):

I'm guessing the equation is \[y= e^{3 ln(x)} \] (it seems too easy with just e^3...) \[ \begin{split} f(x) &= e^{3ln(x)}\\ &= (e^{ln(x)})^{3}\\ &=x^{3}\\ \end{split} \] I think you may want to review some of your Algebra rules for exponents and logarithms! :)

OpenStudy (anonymous):

It's easy either way you look at it.

OpenStudy (accessdenied):

Yeah, true. I guess I should rather say, "its more suspicious as (e^3)l(nx). :P

OpenStudy (anonymous):

i definately need some algebra review. i haven't taken a single math class in 8 years and my school is forcing me to take calculus without allowing me to take the pre-requisites. and i'm trying to graduate this year, so no options but desperately trying to pass

OpenStudy (anonymous):

so, i'm going through my book and doing as many problems as i can and every few hours i go online and ask y'all for help

OpenStudy (anonymous):

thanks so much for putting up with me all afternoon/night

OpenStudy (anonymous):

it's really helped to watch others work out the problems in real time. i've learned a lot from y'all